OpenStudy (anonymous):

1 1 12-- (D) 2--- =? 4 10 5 5--- 6 ---------------------------------------- 1 6--- 40 ---------------------------------------- 2 6--- 5 ---------------------------------------- 1 6--- 4

OpenStudy (ash2326):

\[12\frac{1}{4}\ \div\ 2\frac{1}{10}\] we know that \[a\frac{b}{c}=a+\frac{b}{c}=\frac{ac+b}{c}\] do you get this step?

OpenStudy (anonymous):

no :/

OpenStudy (anonymous):

The answer is C?

OpenStudy (ash2326):

\[a\frac{b}{c}=\text{ This is called mixed fraction}\] for example \[2\frac{1}{6}\] the easy ways of writing this into a fraction is \[2\frac{1}{6}\] \[2+\frac{1}{6}\] next we'll multiply and divide 2 by 6 \[\frac{2\times 6}{6}+\frac{1}{6}\] \[\frac{12}{6}+\frac{1}{6}=\frac{12+1}6=\frac{13}{6}\] Do you understand this?

OpenStudy (anonymous):



OpenStudy (ash2326):

nah, this is not correct

OpenStudy (anonymous):

Oh nvm^^ i know the answer. Thanks for explaining!!!

OpenStudy (ash2326):

how did you find the answer?

OpenStudy (anonymous):

From how you explained it? o.o